# : Timing Interruption Monitoring and Security Assessment for Fronthaul Environments

### v1

@@ 4.1 Attack Methods
Compromised Master Clock. A compromised master clock poses a severe threat to the PTP service, granting attackers control over the network’s synchronization source. Attackers can take control of the master clock by exploiting supply-chain vulnerabilities or gaining unauthorized access, enabling manipulation of timing messages and disruption of network operations.
@@
Within the framework, we focus on two specific attacks targeting PTP synchronization: Spoofing Attack and Replay Attack. These attacks are chosen for their direct impact on PTP synchronization mechanisms. The spoofing attack manipulates the BMCA to influence master clock election, affecting network-wide synchronization. The replay attack disrupts time synchronization by altering drift, offset, and delay calculations on slave clocks. Details of these attacks are illustrated in Fig. 3 and discussed in the following sections. Other attack types, such as DoS and compromised master clock attacks, involve vulnerabilities beyond PTP itself.
4.1.1 Spoofing Attack
The framework spoofing attack targets PTP nodes by sending counterfeit Announce messages to manipulate the BMCA leader election and assume the role of the master clock. Initially, the attacker monitors benign traffic to gather critical information about the protocol configuration and node attributes. With this data, the attacker crafts Announce packets designed to have superior attributes crucial for master clock selection. The BMCA relies on up to nine features to determine the master clock, starting with Priority1 (Sec. 2.2). The attacker manipulates this value by setting it to 1. Similarly, the attacker sets the values for ClockClass to 1. The ClockIdentity field is created by inserting priority values between the manufacturer ID and device ID portions of the MAC address. The attacker inserts ffff ensuring it has higher priority values than the legitimate Master. Other values are copied from the legitimate master.
These crafted messages are sent periodically to maintain the attacker’s role as master. Apart from Announce messages, no synchronization packets are sent, and DelayReq messages from slaves remain unanswered. By disrupting the synchronization process, all clocks operate independently, leading to a gradual drift. Consequently, the attacker introduces synchronization errors among all nodes in the network, causing a complete outage for the gNB, as we will show in Section 6.1 and illustrated in Fig. 1.
4.1.2 Replay Attack
The Replay Attack involves sniffing time synchronization messages (Sync and FollowUp) from the Master clock, storing, and retransmitting them after a delay. When the Slave clock receives the retransmitted messages, it uses the outdated timestamp and a new timestamp to compute drift, delay, and offset (see Fig. 3). The formulas in Sec. 2.2 become: , , and . The actual difference between when is generated and received in the malicious PTP message significantly exceeds the calculated difference, causing offset to increase by orders of magnitude and severely disrupting network synchronization, as we will detail in Section 6.2.
5 Experimental Test Beds
@@
5.1 Production-Ready Network
@@
Our first experimental analysis is performed on a private 5G network, deployed at Northeastern University [43] comprising 8 NVIDIA Aerial RAN CoLab (ARC) nodes, with dedicated Core Network (CN) and FH infrastructure.We utilize this production-ready platform to capture traffic and evaluate attacks on an O-RAN and 5G-compliant system based on a 3GPP 7.2 split, with a combination of open-source components for the higher layers of the stack, as well as commercial devices for the fronthaul infrastructure and the radios [3].
@@
NVIDIA ARC combines the open-source project OpenAirInterface (OAI) [44] for the higher layers of the protocol stack with NVIDIA Aerial, a physical layer implementation accelerated on Graphics Processing Unit (GPU) (NVIDIA A100) and with a 7.2 implementation that combines NVIDIA Mellanox smart Network Interface Cards (NICs). In an ARC server, the GPU is coupled with the programmable NIC (a Mellanox ConnectX-6 Dx) via Remote Direct Memory Access (RDMA), bypassing the CPU for direct packet transfer from the NIC to the GPU. This architecture enables high-speed DU-side termination of the Open FH interface.
@@
The FH infrastructure uses a Dell S5248F-ON switch and a Qulsar QG-2 as the grandmaster clock, distributing PTP and SyncE synchronization to the DU and RU with an O-RAN LLS-C3 configuration. For PTP synchronization, we use ptp4l, an IEEE 1588 compliant implementation for Linux that offers extensive configuration options and enables monitoring and logging of synchronization. The RU is a Foxconn 4T4R unit operating in the GHz band, with Commercial Off-the-Shelf (COTS) 5G smartphones from OnePlus (AC Nord 2003) as User Equipments (UEs) [45].
@@
While the DU codebase is open and potentially extensible to secure the S-Plane, the RU incorporates a closed-source FPGA-based termination for the FH interface, precluding additional security mechanisms in the FH environment from being directly enabled. Thus, we also adopt a trace-based approach similar to [13], configuring a port of the FH switch to mirror the FH traffic to a server running packet capture. Specifically, we capture all traffic traversing the FH for over 20 minutes under several different traffic loads.
5.2 Digital Twin Environment
@@
To leverage these traces we built a digital twin environment using two desktop computers with Intel i9-13900K CPUs and NVIDIA Mellanox ConnectX-4 Lx NICs to represent the DU and RU, while a third desktop computer with an Intel Core i7-10700 CPU serves as the attacker. All machines are connected over a private, 10 Gbps capable switched network. Digital twins provide a highly accurate representation of the network and its devices, making them invaluable for designing and testing ML-based algorithms to detect malicious traffic and ongoing attacks [46]. Once trained, ML-based control solutions need to be validated and tested in controlled environments to avoid disruptions in production networks [47]. Our setup allows for rapid testing, prototyping, and the development of both attacks and solutions, ensuring robust security measures before deployment in real-world scenarios, meeting the need to test and evaluate any proposed S-plane security measures as discussed in [15]. We use .pcap files captured on the production-ready Open FH to accurately emulate the original FH and observe the impact the C and U plane traffic has on the PTP protocol operating in the S-Plane [13].
@@
On top of the background traffic, we start PTP traffic between the DU and RU with ptp4l. This mirrors the setup in our production-ready network, which also uses ptp4l for PTP synchronization. We use the LLS-C2 profile in our digital twin environment, where the DU is the Master Clock and the RU is our Slave Clock. The attacker does not actively participate in the PTP protocol. Instead, it is capable of observing, replaying, or inserting PTP traffic. We discussed the details of our attacks in Section 4.1.
6 Impact of the Attacks
@@
Implementing security measures incurs performance costs [48, 24, 13], so balancing attack risks with the cost of securing the system is crucial. We use the framework to evaluate the impact of attacks on the S-plane within our test environments, following the O-RAN interfaces S-plane specification [20]. Our assessment focuses on the gNB ’s availability. According to NIST’s National Vulnerability Database (NVD)[49], impact levels are categorized as none, low, or high. Low impact attacks cause performance degradation or intermittent availability without fully denying service, affecting functionalities like Carrier Aggregation and handovers[14]. High impact attacks, however, result in complete outages, denying service to all UEs.
6.1 Production-ready Network Impact
@@
We test our attacks in the production-ready private 5G network to assess each attack’s impact following the recently published guidelines in the O-RAN ALLIANCE testing specifications [20]. Based on our threat model in Section 4, we assume that one of the 8 DUs is the compromised machine.
@@
In our initial tests, we found that switch configurations can provide protection against attacks. Malicious packets are not always received by other nodes, as verified through Wireshark captures on both the malicious machine and other nodes. This protection is tied to the switch port’s PTP role settings. On the Dell S5248F-ON switch, ports set to master only send PTP packets, slave mode only receives, and dynamic mode allows both. Additionally, switch configurations can block duplicate MAC addresses. While a properly configured switch prevented our replay attack, assuming correct configurations is risky, especially when DU and RU vendors lack control over network settings, and the complexity and diversity of equipment and configurations in the FH path exacerbate this risk [39, 38, 50].
@@
In our first successful attack, we configure the port of our malicious machine with a PTP master role. We start our private 5G network, connect the UE, and begin sending around 300 Mbps downlink iPerf traffic. After approximately 60 seconds, we launch the spoofing attack, as shown in Fig. 4. Monitoring the PTP service logs at the DU reveals that the attack causes the DU to lose synchronization with the grandmaster clock, leading to a drift in network synchronization as the DU uses its internal clock. This drift degrades UE performance, eventually causing disconnection from the network. About 380 seconds after the beginning of the attack, there is a sudden 50% drop in throughput. From this point, the throughput continues to drop roughly linearly as the clock drift increases. Finally, the DU loses synchronization with the RU, the throughput drops to 0, and the UE disconnects. When this attack stops, the DU resynchronizes with the grandmaster clock, and the network returns to a healthy state.
@@
Next, we configure all ports as PTP dynamic so that malicious packets can be received by the other nodes and repeat the experiment. This time, the malicious PTP packets reach all the machines in the network. Approximately 2 seconds after the attack begins, the RU crashes, stopping its operations, causing the 5G cell to drop and the UE to lose connection, as shown in Fig. 5. While the DU was able to recover on its own after the first attack, the RU required a manual reboot to regain functionality after the second attack. This behavior might be device specific, in our case the Foxconn RU. Regardless of any potential vendor-specific variations, these attacks proved to be very powerful, causing significant disruption and necessitating manual intervention to restore full network operations. These tests demonstrate the impact to the Open RAN gNB is high. Failing to secure the S-plane against spoofing attacks results in a complete outage of the gNB.
6.2 Digital Twin Network Impact
@@
Next we tested the impact of our attacks on PTP functionalities in the Digital Twin environment, focusing on disruptions to synchronization and leader election rather than effects on the UE. We employed cyclic attack/recovery patterns with durations of 30/30, 40/20, 50/10 seconds, and continuous attacks for both of the attacks discussed in Sec. 4.1. The attack duration is the period (in seconds) the attack is carried out, while during the recovery time no attack occurs allowing PTP synchronization to recover. Each experiment involved four background traffic traces, resulting in 32 tests with durations ranging from 120 to 450 seconds, depending on the traffic. Results from the RU confirm significant performance degradation in PTP, with varying severity of impact. Here we briefly highlight the effectiveness of both attacks.
@@
Spoofing Attack: Fig. 6 shows the impact of the 50/10 (50 s attack, 10 s recovery) spoofing attack. Synchronization is disrupted whenever the attacker becomes the master clock, causing the RU to rely on its internal clocks, leading to synchronization drift. The impact of this attack depends on the accuracy of each devices internal clock as well as the length of time of the attack. The observable spike in delay is actually reported shortly after the attack ends, when the legitimate master and slave begin synchronizing again. This spike shows the amount of synchronization drift that occurred during the attack. While there is some natural variation in delay due to the background traffic, shown in light green in Fig. 6, the announce attack (dark red) shows synchronization drift can exceed 40,000 ns.
@@
Replay Attack: This attack has a more severe impact on node synchronization, as shown in Fig. 7, where we display the results of the 30/30 s cycle. Normally, delay values range between 25 and 45 nanoseconds, but during the Replay attack, these values spike by several orders of magnitude, peaking at over 20 million nanoseconds. Comparing Fig. 6 and Fig. 7 it is evident that the replay attack is more disruptive than the spoofing attack, causing a high synchronization drift more quickly and requiring more time for PTP to fully recover.
7 Need for ML Based Detection
@@
As shown in Section 6, synchronization attacks can have catastrophic consequences, including complete outages of the gNB. However, the virtualized, multi-vendor nature of Open RAN and the high precision required for synchronization over the S-plane create challenges for implementing traditional security measures. While hardware-based solutions like MACsec can provide security, they introduce additional costs and complexity to the resource-constrained RU [28, 29].
@@
Rule-Based Solutions: Developing rule-based solutions requires extensive expert knowledge to create specific rules for every potential attack. In Sec. 7.1, we outline the complexity involved in designing expert rules to detect even just two types of attacks.
Fixed Thresholds: Rule-based solutions often rely on fixed thresholds, which are challenging to determine optimally. Even with well-designed thresholds, these methods are limited in their ability to adapt to new or evolving threats. In contrast, ML models do not require setting fixed thresholds. As we will show in Sec. 9.2, well designed models exhibit resilience againste reshaped attacks.
Flexibility: In contrast to rule-based approaches, a well-trained ML model can generalize across various attack types. For instance, we created a new DoS attack by sending a high rate of random or invalid PTP messages. Although this attack did not disrupt PTP timing, it still indicated malicious activity. Our ML model, despite not being specifically trained on this attack, correctly identified over 66% of the malicious traffic, whereas the heuristic model failed completely. This shows that ’s ML-based solution is highly flexible and capable of detecting sophisticated, unpredictable attacks.
Accuracy. An effective monitoring solution must deliver accurate results. As we will demonstrate in Sec. 9.1, our ML model significantly outperforms the rule-based approach, achieving 99% accuracy in a production environment.
Rule-based systems such as those described in [25] struggle to keep pace with the wide range of attacks, while other heuristic methods require additional clocks or changes to the physical infrastructure [33, 16, 36]. Additionally, legitimate changes in master clocks, changes in network topology, and sequence number resets can often resemble malicious activity, complicating detection efforts. ML models excel at identifying patterns and anomalies within large datasets, enabling them to distinguish between benign fluctuations and actual threats. By leveraging ML, we can develop robust monitoring systems capable of adapting to new and evolving attack vectors, ensuring reliable network performance and security.
7.1 Heuristic Detection Results

### v2

@@ 4.1. Attack Methods
Compromised Master Clock. A compromised master clock poses a severe threat to the PTP service, granting attackers control over the networkβs synchronization source. Attackers can take control of the master clock by exploiting supply-chain vulnerabilities or gaining unauthorized access, enabling manipulation of timing messages and disruption of network operations.
@@
Within the framework, we focus on two specific attacks targeting PTP synchronization: Spoofing Attack and Replay Attack. These attacks are chosen for their direct impact on PTP synchronization mechanisms. The spoofing attack manipulates the BMCA to influence master clock election, affecting network-wide synchronization. The replay attack disrupts time synchronization by altering drift, offset, and delay calculations on slave clocks. Details of these attacks are illustrated in Fig.Β 3 and discussed in the following sections. Other attack types, such as DoS and compromised master clock attacks, involve vulnerabilities beyond PTP itself.
4.1.1. Spoofing Attack
The framework spoofing attack targets PTP nodes by sending counterfeit Announce messages to manipulate the BMCA leader election and assume the role of the master clock. Initially, the attacker monitors benign traffic to gather critical information about the protocol configuration and node attributes. With this data, the attacker crafts Announce packets designed to have superior attributes crucial for master clock selection. The BMCA relies on up to nine features to determine the master clock, starting with Priority1 (Sec.Β 2.2). The attacker manipulates this value by setting it to 1. Similarly, the attacker sets the values for ClockClass to 1. The ClockIdentity field is created by inserting priority values between the manufacturer ID and device ID portions of the MAC address. The attacker inserts ffff ensuring it has higher priority values than the legitimate Master. Other values are copied from the legitimate master.
These crafted messages are sent periodically to maintain the attackerβs role as master. Apart from Announce messages, no synchronization packets are sent, and DelayReq messages from slaves remain unanswered. By disrupting the synchronization process, all clocks operate independently, leading to a gradual drift. Consequently, the attacker introduces synchronization errors among all nodes in the network, causing a complete outage for the gNB, as we will show in SectionΒ 6.1 and illustrated in Fig.Β 1.
4.1.2. Replay Attack
The Replay Attack involves sniffing time synchronization messages (Sync and FollowUp) from the Master clock, storing, and retransmitting them after a delay. When the Slave clock receives the retransmitted messages, it uses the outdated timestamp and a new timestamp to compute drift, delay, and offset (see Fig.Β 3). The formulas in Sec.Β 2.2 become: , , and . The actual difference between when is generated and received in the malicious PTP message significantly exceeds the calculated difference, causing offset to increase by orders of magnitude and severely disrupting network synchronization, as we will detail in SectionΒ 6.2.
5. Experimental Test Beds
@@
5.1. Production-Ready Network
@@
Our first experimental analysis is performed on a private 5G network, deployed at Northeastern UniversityΒ (villa2023open) comprising 8Β NVIDIA Aerial RAN CoLab (ARC) nodes, with dedicated Core Network (CN) and FH infrastructure. We utilize this production-ready platform to capture traffic and evaluate attacks on an O-RAN and 5G-compliant system based on a 3GPP 7.2 split, with a combination of open-source components for the higher layers of the stack, as well as commercial devices for the fronthaul infrastructure and the radiosΒ (OranWG4).
@@
NVIDIA ARC combines the open-source project OpenAirInterface (OAI)Β (oai) for the higher layers of the protocol stack with NVIDIA Aerial, a physical layer implementation accelerated on Graphics Processing Unit (GPU) (NVIDIA A100) and with a 7.2 implementation that combines NVIDIA Mellanox smart Network Interface Cards (NICs). In an ARC server, the GPU is coupled with the programmable NIC (a Mellanox ConnectX-6 Dx) via Remote Direct Memory Access (RDMA), bypassing the CPU for direct packet transfer from the NIC to the GPU. This architecture enables high-speed DU-side termination of the Open FH interface.
@@
The FH infrastructure uses a Dell S5248F-ON switch and a Qulsar QG-2 as the grandmaster clock, distributing PTP and SyncE synchronization to the DU and RU with an O-RAN LLS-C3 configuration. For PTP synchronization, we use ptp4l, an IEEE 1588 compliant implementation for Linux that offers extensive configuration options and enables monitoring and logging of synchronization. The RU is a Foxconn 4T4R unit operating in the GHz band, with Commercial Off-the-Shelf (COTS) 5G smartphones from OnePlus (AC Nord 2003) as User Equipments (UEs)Β (oneplus).
@@
While the DU codebase is open and potentially extensible to secure the S-Plane, the RU incorporates a closed-source FPGA-based termination for the FH interface, precluding additional security mechanisms in the FH environment from being directly enabled. Thus, we also adopt a trace-based approach similar toΒ (groen2024securing), configuring a port of the FH switch to mirror the FH traffic to a server running packet capture. Specifically, we capture all traffic traversing the FH for over 20 minutes under several different traffic loads.
5.2. Digital Twin Environment
@@
To leverage these traces we built a digital twin environment using two desktop computers with Intel i9-13900K CPUs and NVIDIA Mellanox ConnectX-4 Lx NICs to represent the DU and RU, while a third desktop computer with an Intel Core i7-10700 CPU serves as the attacker. All machines are connected over a private, 10 Gbps capable switched network. Digital twins provide a highly accurate representation of the network and its devices, making them invaluable for designing and testing ML-based algorithms to detect malicious traffic and ongoing attacksΒ (10179151). Once trained, ML-based control solutions need to be validated and tested in controlled environments to avoid disruptions in production networksΒ (polese2024colosseum). Our setup allows for rapid testing, prototyping, and the development of both attacks and solutions, ensuring robust security measures before deployment in real-world scenarios, meeting the need to test and evaluate any proposed S-plane security measures as discussed in (s_plane_survey). We use .pcap files captured on the production-ready Open FH to accurately emulate the original FH and observe the impact the C and U plane traffic has on the PTP protocol operating in the S-PlaneΒ (groen2024securing).
@@
On top of the background traffic, we start PTP traffic between the DU and RU with ptp4l. This mirrors the setup in our production-ready network, which also uses ptp4l for PTP synchronization. We use the LLS-C2 profile in our digital twin environment, where the DU is the Master Clock and the RU is our Slave Clock. The attacker does not actively participate in the PTP protocol. Instead, it is capable of observing, replaying, or inserting PTP traffic. We discussed the details of our attacks in SectionΒ 4.1.
6. Impact of the Attacks
@@
Implementing security measures incurs performance costsΒ (groen2023implementing; groen2024securing), so balancing attack risks with the cost of securing the system is crucial. We use the framework to evaluate the impact of attacks on the S-plane within our test environments, following the O-RAN interfaces S-plane specificationΒ (OranWG11-sectestspec). Our assessment focuses on the gNB βs availability. According to NISTβs National Vulnerability Database (NVD)(NVD), impact levels are categorized as none, low, or high. Low impact attacks cause performance degradation or intermittent availability without fully denying service, affecting functionalities like Carrier Aggregation and handovers(municio2023ran). High impact attacks, however, result in complete outages, denying service to all UEs.
6.1. Production-ready Network Impact
@@
We test our attacks in the production-ready private 5G network to assess each attackβs impact following the recently published guidelines in the O-RAN ALLIANCE testing specificationsΒ (OranWG11-sectestspec). Based on our threat model in SectionΒ 4, we assume that one of the 8Β DUs is the compromised machine.
@@
In our initial tests, we found that switch configurations can provide protection against attacks. Malicious packets are not always received by other nodes, as verified through Wireshark captures on both the malicious machine and other nodes. This protection is tied to the switch portβs PTP role settings. On the Dell S5248F-ON switch, ports set to master only send PTP packets, slave mode only receives, and dynamic mode allows both. Additionally, switch configurations can block duplicate MAC addresses. While a properly configured switch prevented our replay attack, assuming correct configurations is risky, especially when DU and RU vendors lack control over network settings, and the complexity and diversity of equipment and configurations in the FH path exacerbate this riskΒ (verizon; ibm2023; rapid7).
@@
In our first successful attack, we configure the port of our malicious machine with a PTP master role. We start our private 5G network, connect the UE, and begin sending around 300 Mbps downlink iPerf traffic. After approximately 60 seconds, we launch the spoofing attack, as shown in Fig.Β 4. Monitoring the PTP service logs at the DU reveals that the attack causes the DU to lose synchronization with the grandmaster clock, leading to a drift in network synchronization as the DU uses its internal clock. This drift degrades UE performance, eventually causing disconnection from the network. About 380 seconds after the beginning of the attack, there is a sudden 50% drop in throughput. From this point, the throughput continues to drop roughly linearly as the clock drift increases. Finally, the DU loses synchronization with the RU, the throughput drops to 0, and the UE disconnects. When this attack stops, the DU resynchronizes with the grandmaster clock, and the network returns to a healthy state.
@@
Next, we configure all ports as PTP dynamic so that malicious packets can be received by the other nodes and repeat the experiment. This time, the malicious PTP packets reach all the machines in the network. Approximately 2 seconds after the attack begins, the RU crashes, stopping its operations, causing the 5G cell to drop and the UE to lose connection, as shown in Fig.Β 5. While the DU was able to recover on its own after the first attack, the RU required a manual reboot to regain functionality after the second attack. This behavior might be device specific, in our case the Foxconn RU. Regardless of any potential vendor-specific variations, these attacks proved to be very powerful, causing significant disruption and necessitating manual intervention to restore full network operations. These tests demonstrate the impact to the Open RAN gNB is high. Failing to secure the S-plane against spoofing attacks results in a complete outage of the gNB.
6.2. Digital Twin Network Impact
@@
Next we tested the impact of our attacks on PTP functionalities in the Digital Twin environment, focusing on disruptions to synchronization and leader election rather than effects on the UE. We employed cyclic attack/recovery patterns with durations of 30/30, 40/20, 50/10 seconds, and continuous attacks for both of the attacks discussed in Sec.Β 4.1. The attack duration is the period (in seconds) the attack is carried out, while during the recovery time no attack occurs allowing PTP synchronization to recover. Each experiment involved four background traffic traces, resulting in 32 tests with durations ranging from 120 to 450 seconds, depending on the traffic. Results from the RU confirm significant performance degradation in PTP, with varying severity of impact.
To generate labeled training data, we used logs to identify the timing of attacks and the specific machines involved (based on MAC addresses). A script was written to label each packet in the dataset as either benign (0) or malicious (1) based on whether it came from an attacking machine during the attack window.
Comprehensive results are displayed in the Appendix (Sec.Β A.1), but here we highlight the effectiveness of both attacks.
@@
Spoofing Attack: Fig.Β 6 shows the impact of the 50/10 (50Β s attack, 10Β s recovery) spoofing attack. Synchronization is disrupted whenever the attacker becomes the master clock, causing the RU to rely on its internal clocks, leading to synchronization drift. The impact of this attack depends on the accuracy of each devices internal clock as well as the length of time of the attack. The observable spike in delay is actually reported shortly after the attack ends, when the legitimate master and slave begin synchronizing again. This spike shows the amount of synchronization drift that occurred during the attack. While there is some natural variation in delay due to the background traffic, shown in light green in Fig.Β 6, the announce attack (dark red) shows synchronization drift can exceed 40,000 ns.
@@
Replay Attack: This attack has a more severe impact on node synchronization, as shown in Fig.Β 7, where we display the results of the 30/30Β s cycle. Normally, delay values range between 25 and 45 nanoseconds, but during the Replay attack, these values spike by several orders of magnitude, peaking at over 20 million nanoseconds. Comparing Fig.Β 6 and Fig.Β 7 it is evident that the replay attack is more disruptive than the spoofing attack, causing a high synchronization drift more quickly and requiring more time for PTP to fully recover.
7. Need for ML Based Detection
@@
As shown in SectionΒ 6, synchronization attacks can have catastrophic consequences, including complete outages of the gNB. However, the virtualized, multi-vendor nature of Open RAN and the high precision required for synchronization over the S-plane create challenges for implementing traditional security measures. While hardware-based solutions like MACsec can provide security, they introduce additional costs and complexity to the resource-constrained RUΒ (dik2023macsec; dik2023100).
@@
Rule-Based Solutions: Developing rule-based solutions requires extensive expert knowledge to create specific rules for every potential attack. In Sec.Β 7.1, we outline the complexity involved in designing expert rules to detect even just two types of attacks.
Fixed Thresholds: Rule-based solutions often rely on fixed thresholds, which are challenging to determine optimally. Even with well-designed thresholds, these methods are limited in their ability to adapt to new or evolving threats. In contrast, ML models do not require setting fixed thresholds. As we will show in Sec.Β 9.2, well designed models exhibit resilience againste reshaped attacks.
Flexibility: In contrast to rule-based approaches, a well-trained ML model can generalize across various attack types. For instance, we created a new DoS attack by sending a high rate of random or invalid PTP messages. Although this attack did not disrupt PTP timing, it still indicated malicious activity. Our ML model, despite not being specifically trained on this attack, correctly identified over 66% of the malicious traffic, whereas the heuristic model failed completely. This shows that βs ML-based solution is highly flexible and capable of detecting sophisticated, unpredictable attacks.
Accuracy. An effective monitoring solution must deliver accurate results. As we will demonstrate in Sec.Β 9.1, our ML model significantly outperforms the rule-based approach, achieving 99% accuracy in a production environment.
Rule-based systems such as those described in (ptp_standard) struggle to keep pace with the wide range of attacks, while other heuristic methods require additional clocks or changes to the physical infrastructure (moussa2016detection; shi2023ms; finkenzeller2024ptpsec). Additionally, legitimate changes in master clocks, changes in network topology, and sequence number resets can often resemble malicious activity, complicating detection efforts. ML models excel at identifying patterns and anomalies within large datasets, enabling them to distinguish between benign fluctuations and actual threats. By leveraging ML, we can develop robust monitoring systems capable of adapting to new and evolving attack vectors, ensuring reliable network performance and security.
7.1. Heuristic Detection Results